OpenStudy (anonymous):

Simplify the following exponential expressions. Explain how. =) (-5ab)^3 (-8)^0 + 2^0 (x^5)(x^2)( x^3) (y4)^9 -6^2

OpenStudy (anonymous):

1. (-5ab)^3 = -5^3 * a ^ 3 * b ^3 = -15a^3b^3 2. any number ^ 0 always equals 1 so (-8)^0 + 2^0 = 1 + 1 = 2 3. (x^5)(x^2)( x^3) = x ^ (5+2+3) = x^10 4. (y4)^9 = i guess you meant (y^4)^9 = y ^ (9*4) = y ^36 5. -6^2 = 12